# NavThinker: Action-Conditioned World Models for Coupled Prediction and Planning in Social Navigation

### 16 Mar 2026 v1

@@ Abstract
Social navigation requires robots to act safely in dynamic human environments. Effective behavior demands thinking ahead: reasoning about how the scene and pedestrians evolve under different robot actions rather than reacting to current observations alone. This creates a coupled prediction-planning challenge, where robot actions and human motion mutually influence each other. To address this challenge, we propose NavThinker, a future-aware framework that couples an action-conditioned world model with on-policy reinforcement learning. The world model operates in the Depth Anything V2 patch feature space and performs autoregressive prediction of future scene geometry and human motion; multi-head decoders then produce future depth maps and human trajectories, yielding a future-aware state aligned with traversability and interaction risk. Crucially, we train the policy with DD-PPO while injecting world-model think-ahead signals via: (i) action-conditioned future features fused into the current observation embedding and (ii) social reward shaping from predicted human trajectories. Experiments on single- and multi-robot Social-HM3D show state-of-the-art navigation success, with zero-shot transfer to Social-MP3D and real-world deployment on a Unitree Go2, validating generalization and practical applicability. Webpage: https://github.com/hutslib/NavThinker.
@@ I Introduction
Figure 1: NavThinker: future-aware social navigation. Social navigation requires the robot to reach a goal in environments shared with dynamic pedestrians. Red regions highlight interaction zones where the robot and pedestrians may conflict. Top: egocentric depth observations at two timesteps. Bottom: the world model imagines action-conditioned futures for the corresponding observations. By imagining how the scene evolves under different candidate actions, NavThinker anticipates potential social conflicts and selects safe, socially compliant actions.
@@
Figure 2: Overview of the NavThinker framework. Our framework consists of two modules: a world model that learns action-conditioned scene dynamics (top), and an imagination-augmented planner policy trained with DD-PPO (bottom). During World Model Learning, depth observations, actions, and robot states are extracted from the Replay Buffer. A frozen DA-V2 ViT encoder extracts patch embeddings, and a Causal Attention Transformer autoregressively predicts future latent features. A Depth Decoder and a Human Trajectory Decoder are trained alongside a latent consistency loss to anchor representations to scene geometry and human motion. During Policy Learning, the Observation Encoder produces the current embedding from depth and robot states. The Imagination module queries the world model under each candidate action, generating action-conditioned future features that are fused with the current embedding via Feature Fusion. The fused representation feeds into the $D$-PPO Actor-Critic network. Predicted human trajectories additionally provide Reward Shaping, coupling prediction with planning in an imagine-then-act loop.
@@ III Problem Formulation
We consider a point-goal SocialNav task in environments populated by $N$ dynamic humans. Due to egocentric sensing, occlusions, and limited field-of-view, the environment state can only be partially observed. We formulate this task as a Partially Observable Markov Decision Process (POMDP) $\left(\mathcal{S},\mathcal{O},\mathcal{A},\mathcal{T},r,\gamma\right)$, where $s_{t}\in\mathcal{S}$ is the latent environment state, $o_{t}\in\mathcal{O}$ is the observation, $a_{t}\in\mathcal{A}$ is the action, $\mathcal{T}(s_{t+1}\!\mid s_{t},a_{t})$ is the transition kernel, $r(s_{t},a_{t})$ is the reward, and $\gamma\in(0,1)$ is the discount factor. At each timestep $t$, the agent selects $a_{t}$ based on the observation history $o_{\leq t}$.
@@ IV-A Overview
NavThinker tackles coupled prediction and planning in social navigation by learning a world model and using it as an imagination engine for policy improvement. As shown in Fig. 2, it consists of two modules: (i) an action-conditioned scene-and-interaction world model operating in DA-V2 patch-feature space, which predicts future patch latents and uses multi-head decoders to forecast future depth and human trajectories, producing a future-aware state capturing traversability and interaction risk; and (ii) an imagination-augmented planner policy trained on-policy with DD-PPO, which exploits imagined rollouts by fusing candidate-action-conditioned future features with the current observation embedding and by shaping rewards from predicted human trajectories. This imagination-enabled training couples prediction and planning through the learned latent dynamics.
@@ IV-C Imagination-Augmented Policy Learning
We train the policy with DD-PPO [35] under a reward that combines task progress with world-model-derived social shaping:

### 18 Mar 2026 v2

@@ Abstract
Social navigation requires robots to act safely in dynamic human environments. Effective behavior demands thinking ahead: reasoning about how the scene and pedestrians evolve under different robot actions rather than reacting to current observations alone. This creates a coupled prediction-planning challenge, where robot actions and human motion mutually influence each other. To address this challenge, we propose NavThinker, a future-aware framework that couples an action-conditioned world model with on-policy reinforcement learning. The world model operates in the Depth Anything V2 patch feature space and performs autoregressive prediction of future scene geometry and human motion; multi-head decoders then produce future depth maps and human trajectories, yielding a future-aware state aligned with traversability and interaction risk. Crucially, we train the policy with $D$-PPO while injecting world-model think-ahead signals via: (i) action-conditioned future features fused into the current observation embedding and (ii) social reward shaping from predicted human trajectories. Experiments on single- and multi-robot Social-HM3D show state-of-the-art navigation success, with zero-shot transfer to Social-MP3D and real-world deployment on a Unitree Go2, validating generalization and practical applicability. Webpage: https://hutslib.github.io/NavThinker/.
@@ I Introduction
Figure 1: NavThinker: future-aware social navigation. Social navigation requires the robot to reach a goal in environments shared with dynamic pedestrians. Red regions highlight interaction zones where the robot and pedestrians may conflict. Top: egocentric depth observations. Bottom: the world model imagines action-conditioned futures for the corresponding observations. By imagining how the scene evolves under different candidate actions, NavThinker anticipates potential social conflicts and selects safe, socially compliant actions.
@@
Figure 2: Overview of the NavThinker framework. Our framework consists of two modules: a world model that learns action-conditioned scene dynamics (top), and an imagination-augmented planner policy trained with $D$-PPO (bottom). During World Model Learning, depth observations, actions, and robot states are extracted from the Replay Buffer. A frozen DA-V2 ViT encoder extracts patch embeddings, and a Causal Attention Transformer autoregressively predicts future latent features. A Depth Decoder and a Human Trajectory Decoder are trained alongside a latent consistency loss to anchor representations to scene geometry and human motion. During Policy Learning, the Observation Encoder produces the current embedding from depth and robot states. The Imagination module queries the world model under each candidate action, generating action-conditioned future features that are fused with the current embedding via Feature Fusion. The fused representation feeds into the $D$-PPO Actor-Critic network. Predicted human trajectories additionally provide Reward Shaping, coupling prediction with planning in an imagine-then-act loop.
@@ III Problem Formulation
We consider a point-goal SocialNav task in environments populated by NN dynamic humans. Due to egocentric sensing, occlusions, and limited field-of-view, the environment state can only be partially observed. We formulate this task as a Partially Observable Markov Decision Process (POMDP) $\left(\mathcal{S},\mathcal{O},\mathcal{A},\mathcal{T},r,\gamma\right)$, where $s_{t}\in\mathcal{S}$ is the latent environment state, $o_{t}\in\mathcal{O}$ is the observation, $a_{t}\in\mathcal{A}$ is the action, $\mathcal{T}(s_{t+1}\!\mid s_{t},a_{t})$ is the transition kernel, $r(s_{t},a_{t})$ is the reward, and $\gamma\in(0,1)$ is the discount factor. At each timestep $t$, the agent selects $a_{t}$ based on the observation history $o_{\leq t}$.
@@ IV-A Overview
NavThinker tackles coupled prediction and planning in social navigation by learning a world model and using it as an imagination engine for policy improvement. As shown in Fig. 2, it consists of two modules: (i) an action-conditioned scene-and-interaction world model operating in DA-V2 patch-feature space, which predicts future patch latents and uses multi-head decoders to forecast future depth and human trajectories, producing a future-aware state capturing traversability and interaction risk; and (ii) an imagination-augmented planner policy trained on-policy with $D$-PPO, which exploits imagined rollouts by fusing candidate-action-conditioned future features with the current observation embedding and by shaping rewards from predicted human trajectories. This imagination-enabled training couples prediction and planning through the learned latent dynamics.
@@ IV-C Imagination-Augmented Policy Learning
We train the policy with $D$-PPO [35] under a reward that combines task progress with world-model-derived social shaping: